A motorboat can travel 28 miles downstream in the same amount of time as it can travel 18 miles upstream. Find the speed of the current if the boat can travel 23 mph in still water. I thought the problem should be something like 28/23+x=18/23-x but this is not giving me the correct answer...not sure what else I can do
Answer by Alan3354(69443)

A motorboat can travel 28 miles downstream in the same amount of time as it can travel 18 miles upstream. Find the speed of the current if the boat can travel 23 mph in still water. I thought the problem should be something like 28/23+x=18/23-x but this is not giving me the correct answer...not sure what else I can do
---------------
28/(23+x) = 18/(23-x) looks right to me
Cross multiply
28*(23-x) = 18*(23+x)
644 - 28x = 414 + 18x
230 = 46x
x = 5 mph
-------